the distance between the two slits in a double-slit experiment is 0.040mm. the second-order bright fringe (m=2) is measured at a distance of 3.5 cm from the central maximum on a screen placed 0.90m from the slits. what is the wavelength of the light?
Answer #1

yn = n*wavelength*D/d

3.5e-2 = (2*wavelength *0.9)/0.04e-3)

wavelength = 3.5e-2*0.04e-3/0.18 = 7.78*10^-6 m <-answer